OpenStudy (anonymous):

give the equation of the line that goes through the points (2, -1) and (5,2)

jimthompson5910 (jim_thompson5910):

first you need the slope of the line that goes through the two points so use the slope formula m = (y2-y1)/(x2-x1)

jimthompson5910 (jim_thompson5910):

what do you get for the slope?

OpenStudy (anonymous):

1...correct?

jimthompson5910 (jim_thompson5910):

correct, m = 1

jimthompson5910 (jim_thompson5910):

now use y = mx+b

jimthompson5910 (jim_thompson5910):

y = mx + b y = 1*x + b ... plug in m = 1 -1 = 1*(2) + b ... plug in x = 2 and y = -1, from the first point -1 = 2 + b ... ... ... b = ???

OpenStudy (anonymous):

is it 2??

jimthompson5910 (jim_thompson5910):

-1 = 2 + b -1 - 2 = 2 + b - 2 -3 = b b = -3

jimthompson5910 (jim_thompson5910):

m = 1 b = -3 y = mx+b turns into y = 1x + (-3) and that simplifies to y = x - 3